# LLMLight: Large Language Models as Traffic Signal Control Agents

### v4

@@ Abstract.
Traffic Signal Control (TSC) is a crucial component in urban traffic management, aiming to optimize road network efficiency and reduce congestion. Traditional methods in TSC, primarily based on transportation engineering and reinforcement learning (RL), often exhibit limitations in generalization across varied traffic scenarios and lack interpretability. This paper presents LLMLight, a novel framework employing Large Language Models (LLMs) as decision-making agents for TSC. Specifically, the framework begins by instructing the LLM with a knowledgeable prompt detailing real-time traffic conditions. Leveraging the advanced generalization capabilities of LLMs, LLMLight engages a reasoning and decision-making process akin to human intuition for effective traffic control. Moreover, we build LightGPT, a specialized backbone LLM tailored for TSC tasks. By learning nuanced traffic patterns and control strategies, LightGPT enhances the LLMLight framework cost-effectively. Extensive experiments on nine real-world and synthetic datasets showcase the remarkable effectiveness, generalization ability, and interpretability of LLMLight against nine transportation-based and RL-based baselines.
@@
1. Introduction
Traffic congestion has emerged as a critical issue impacting both human society and the environment. As urban migration continues to accelerate, city populations are expanding rapidly, further intensifying this challenge. This trend has necessitated a focused effort on optimizing Traffic Signal Control (TSC), a pivotal research area in intelligent transportation (Wu et al., 2023; Zhang et al., 2023). Efficient TSC not only promises substantial economic and environmental benefits but also enhances societal well-being. However, effectively managing traffic signals in complex and dynamic urban road networks remains a daunting task.
Existing research on TSC has primarily fallen into two categories: transportation (Koonce and Rodegerdts, 2008; Hunt et al., 1982; Lowrie, 1990; Varaiya, 2013; Cools et al., 2013) and Reinforcement Learning (RL)-based approaches (Chen et al., 2020; Wei et al., 2019b; Oroojlooy et al., 2020; Chenguang et al., 2021; Wu et al., 2021; Zhang et al., 2022b; Wei et al., 2019a; Wu et al., 2023). Transportation methods primarily focus on crafting efficient heuristic algorithms, dynamically adapting traffic signal configurations based on lane-level traffic conditions. However, these methods heavily rely on manual design, demanding substantial human effort. The emergence of deep neural networks (DNNs) (Mnih et al., 2015) led to the introduction of deep RL-based techniques to address this challenge (Abdulhai et al., 2003; El-Tantawy and Abdulhai, 2010). These approaches have exhibited remarkable performance across various traffic scenarios. Nevertheless, RL-based methods also present several drawbacks. Primarily, they may struggle with limited generalization ability, particularly when transferring to larger-scale road networks or under highly uncommon scenarios (e.g., extreme high-traffic situations), as their training data only covers limited traffic situations. Additionally, RL-based methods lack interpretability due to the black-box nature of DNNs, which makes it hard to explain the rationale behind their control actions under specific traffic conditions.
The recent emergence of Large Language Models (LLMs) has revolutionized several domains with their exceptional zero-shot learning and generalization capabilities, emulating human-like reasoning to solve complex tasks. For instance, AutoGPT (et al., 2023) proposes to break tasks into multiple sub-goals and iterate until the main task is completed. In the transportation domain, GPT-Driver (Mao et al., 2023) proposes to instruct GPT-3.5 to tackle motion planning tasks in autonomous driving. For traffic control, PromptGAT (Da et al., 2023) uses LLMs to generate human knowledge to help the DNN model understand long-tail cases (e.g., extreme weather) in TSC tasks, aiming to bridge the gap between real-world and simulations. However, while existing research utilizes LLMs as auxiliary tools to enhance decision-making, the direct employment of LLMs as TSC agents for human-like decision-making remains largely unexplored. In particular, two major challenges arise toward this goal.
First, LLMs are typically pre-trained on large-scale natural language corpora and rarely incorporate non-textual traffic data, such as sensor readings and GPS trajectories (Touvron et al., 2023; Chung et al., 2022). Despite their generalization capability across various tasks and domains, an inherent gap exists between real-time traffic data and linguistic understanding. The first challenge lies in enabling LLMs to comprehend real-time traffic dynamics and effectively interact with the traffic environment, which is critical for effective LLM-based traffic signal control.
@@ 1. Introduction
To this end, we introduce LLMLight, a traffic signal control agent framework based on LLMs, as depicted in Figure 1. Specifically, we consider TSC as a partially observable Markov Game, where each agent, armed with an LLM, manages the traffic light at an intersection. At each signal-switching time step, the agent collects traffic conditions of the target intersection and transforms them into human-readable text as real-time observation. Additionally, we incorporate task descriptions enriched with commonsense knowledge about a control strategy to aid the LLM’s understanding of traffic management tasks. The combined real-time observation, task description, and control action space form a knowledgeable prompt that guides the LLM’s decision-making. Then, the agent leverages Chain-of-Thought (CoT) reasoning to determine the optimal traffic signal configuration for the subsequent time step.
@@
Furthermore, we construct a specialized LLM, LightGPT, to enhance the LLMLight framework. On the one hand, we propose imitation fine-tuning to let the specialized LLM exploit high-quality control actions and underlying rationales derived from GPT-4. On the other hand, we introduce a policy refinement process that utilizes a well-trained critic model to evaluate and improve the agent’s actions. The optimized LightGPT produces more effective control policies and showcases remarkable generalization ability across diverse traffic scenarios in a more cost-effective manner than GPT-4.
@@
We summarize the major contributions of this paper as follows: (1) We pioneer the use of advanced LLMs as control agents, empowering human-like decision-making for interpretable traffic signal control. To our knowledge, this is the first exploration of LLMs as decision-making agents in TSC. (2) We introduce LLMLight [^1] Our project is available at https://github.com/usail-hkust/LLMTSCS, a tailored framework that engages various LLMs in effectively regulating traffic signals via knowledgeable instructions. (3) We construct LightGPT [^2] The model weights is available at https://huggingface.co/USAIL-HKUSTGZ/LLMLight-LightGPT, a specialized LLM backbone, which is optimized for TSC tasks and significantly enhances the framework’s capability to derive effective control policies in a cost-effective manner. (4) Extensive experiments on nine traffic flow datasets demonstrate the superior effectiveness, generalization capabilities, and interpretability of our framework. Additionally, these experiments reveal key insights into the potential of integrating LLMs into intelligent transportation systems.
@@ 2. Preliminary
Please refer to Appendix A.1 (Figure 9) for more details of the most used setting of the intersection, lanes, and signal phases.
@@ 3.2.1. Observation Feature Construction
Subsequently, these features on every incoming lane (where the vehicle enters the intersection) are transformed into human-readable text as the observation $o$ of the LLM.
@@ 3.2.2. Commonsense Augmented Prompt Construction
In addition to observation, we provide the LLM with comprehensive information essential in addressing traffic signal control, including traffic scenario description $d_{scene}$, task descriptions $d_{task}$, and control action space $A$. This enables the LLM to thoroughly understand the task, facilitating reasonable control actions. Furthermore, we integrate commonsense knowledge $d_{know}$ to mitigate generalist LLMs’ limitations in traffic management, a domain they are not explicitly trained for. Specifically, this knowledge dictates that the agent needs to prioritize lanes with extended queue lengths while paying less attention to vehicles distant from the intersection. Formally, we denote the composed prompt as $X=Prompt(o,d_{scene},d_{task},d_{know},A)$. A glimpse of the prompt template is shown in Figure 1(a), and a detailed example is provided in Appendix A.5 (Table 4).
@@ 3.3. Backbone LLM Training
Fine-tuning has become a common routine to align LLM with diverse downstream applications (Chen et al., 2023a; Zeng et al., 2023) by augmenting the domain expertise in the specific realm. We propose a training procedure to optimize an LLM specifically tailored for traffic signal control, leading to the development of LightGPT. Figure 3 presents the training workflow. It primarily consists of three stages: (1) reasoning trajectory collection and filtering: it collects the CoT reasoning trajectory of GPT-4 for imitation fine-tuning, then we select the trajectories that most align with the long-term goal to ensure the data quality; (2) imitation fine-tuning: it leverages the collected reasoning trajectories for training the backbone LLM on actions taken by GPT-4 and their underlying rationales; (3) critic-guided policy refinement: it further improves the decision-making process of the LLM by fine-tuning with the feedback of a well-trained critic model.
@@ 3.3.1. Trajectory Collection and Filtering
where $o_{t}$ and $a_{t}$ are the observation and action chosen at the signal-switching time step $t$, $\gamma\in[0,1]$ is the discount factor. $R(o,a)$ is the reward function that provides feedback for executing action $a$ under observation $o$ (e.g., the negative of queue length). $Q(o,a)$ is the action-value function estimating the cumulative future reward earned after executing $a$. Actions estimated to yield higher future rewards suggest greater alignment with the long-term goal.
Subsequently, the trained action-value function is utilized as a critic to evaluate the decision-making of GPT-4. We selectively retain only reasoning trajectories where the selected actions are associated with the highest estimated future reward, as:
@@ 3.3.2. Imitation Fine-tuning
where $P_{\pi_{\theta}}(y_{w}|X,Y_{<w})$ is the probability of being prompted with $X$ to generate the token $y_{w}$. Due to the enormous GPU memory usage of full model training, we adopt parameter-efficient fine-tuning (PEFT) techniques, specifically low-rank adaptation (LoRA) (Hu et al., 2022), in our training. LoRA only adds small adapters on top of some specific layers, drastically reducing the number of trainable parameters.
@@ 3.3.3. Critic-guided Policy Refinement
To further enhance the effectiveness of the LLM’s control policy, we propose a policy refinement procedure, which adjusts the LLM’s reasoning trajectory for directing better actions. Similarly, we use the pre-trained action-value function elaborated in Section 3.3.1 as the critic model to evaluate the actions derived by the backbone LLM. Higher evaluation scores correspond to greater estimated future rewards. Subsequently, we implement an alignment fine-tuning algorithm to adjust the reasoning trajectory, ultimately guiding the LLM towards actions that yield higher scores. Inspired by studies like (Wang et al., 2023a; Yuan et al., 2023), we leverage a ranking loss function to achieve this goal.
@@
where $p_{j*}=\min\limits_{q_{k}>q_{j}}p_{k}$ is the probability of the least favorable reasoning trajectory rated higher than $Y_{j}$, $\beta$ is a hyper-parameter, $e^{(p_{j}-p_{i})}$ is the alignment term to elevate the probability of generating trajectories associated with actions rated higher scores, and $e^{(2p_{j*}-2\beta-p_{i}-p_{j})}$ is the constraint term for prohibiting performance degradation.
It’s crucial to aware that while this refinement procedure is guided by a pre-trained critic model developed from RL, this doesn’t imply that the backbone LLM learns exclusively from it. The alignment process is trained on reasoning trajectories sampled by the LLM’s policy. The objective is to encourage the model to discern which trajectory is more reasonable and leads to decisions more closely aligned with long-term goals. However, it’s important to note that we do not provide any ground-truth labels, thus the LLM is not bound to follow the critic’s predicted “best” action directly. The experiments presented in Subsection 4.6 validate that RL models experience performance degradation under varied traffic scenarios, whereas LLMLight equipped with LightGPT remains robust based on our proposed training procedure.
@@ 4.1.1. Environment Settings
We conduct experiments on an open-sourced simulator, CityFlow (Zhang et al., 2019), to evaluate the efficiency of compared methods. It provides extensive APIs to access the traffic state features and execute the selected control actions. In our experiments, each intersection connects four lane sections: East, West, North, and South. We utilize four signal phases as corresponding control actions, including ETWT (go-through from east and west), ELWL (left-turn from east and west), NTST (go-through from north and south), and NLSL (left-turn from north and south). The turn-right movements are always allowed in simulation. The green signal phase duration is set to thirty seconds. Following existing studies (Zhang et al., 2022b; Wei et al., 2019b), a three-second yellow signal and a two-second all-red time follows each green signal to prepare the transition. The simulation spans one hour.
4.1.2. Datasets
Our experiments use seven real-world (Wei et al., 2019c) and two synthetic traffic flow datasets. Statistics for these datasets are detailed in Table 1. Visualizations of road networks are presented in Appendix A.4 (Figure 11).
  • Jinan: These datasets are collected from the Dongfeng sub-district, Jinan, China, featuring 12 intersections. The East and West sections span 400 meters, while the South and North sections cover 800 meters. Across different periods, there are three traffic flow datasets recorded in this area.
  • Hangzhou: These datasets are collected in the Gudang sub-district, Hangzhou, China. There is a total of 16 intersections in this area. The East and West sections span 800 meters, while the South and North sections extend 600 meters. Across different periods, this dataset comprises two traffic flow datasets from the recorded area.
  • New York: To evaluate the scalability of different methods, we introduce two large-scale traffic datasets collected in the Upper East Side of Manhattan, New York, USA, utilizing publicly available taxi trip data. It encompasses 196 intersections within the region. All lanes span 300 meters. Over different periods, the dataset includes two distinct traffic flow datasets.
  • Extreme High-traffic Sceanrios: To simulate high-traffic scenarios, our experiments encompass two synthetic traffic flow datasets on the Jinan and Hangzhou road networks, featuring about four times more vehicles arriving within each 5-minute interval than the original datasets.

### v5

@@ Abstract.
Traffic Signal Control (TSC) is a crucial component in urban traffic management, aiming to optimize road network efficiency and reduce congestion. Traditional TSC methods, primarily based on transportation engineering and reinforcement learning (RL), often struggle with generalization abilities across varied traffic scenarios and lack interpretability. This paper presents LLMLight, a novel framework employing Large Language Models (LLMs) as decision-making agents for TSC. Specifically, the framework begins by instructing the LLM with a knowledgeable prompt detailing real-time traffic conditions. Leveraging the advanced generalization capabilities of LLMs, LLMLight engages a reasoning and decision-making process akin to human intuition for effective traffic control. Moreover, we build LightGPT [^1] The model weights are available at https://huggingface.co/lightgpt., a specialized backbone LLM tailored for TSC tasks. By learning nuanced traffic patterns and control strategies, LightGPT enhances the LLMLight framework cost-effectively. Extensive experiments conducted on ten real-world and synthetic datasets, along with evaluations by fifteen human experts, demonstrate the exceptional effectiveness, generalization ability, and interpretability of LLMLight with LightGPT, outperforming nine baseline methods and ten advanced LLMs. Our project is available at https://github.com/usail-hkust/LLMTSCS.
@@
1. Introduction
Traffic congestion has emerged as a critical issue impacting both human society and the environment. As urban migration continues accelerating, city populations are expanding rapidly, further intensifying this challenge. This trend has necessitated a focused effort on optimizing Traffic Signal Control (TSC), a pivotal research area in intelligent transportation (Wu et al., 2023; Zhang et al., 2023b). Efficient TSC not only promises substantial economic and environmental benefits but also enhances societal well-being. However, effectively managing traffic signals in complex and dynamic urban road networks remains daunting.
Existing work on TSC primarily falls into two categories: transportation (Koonce and Rodegerdts, 2008; Hunt et al., 1982; Lowrie, 1990; Varaiya, 2013; Cools et al., 2013) and Reinforcement Learning (RL)-based approaches (Chen et al., 2020; Wei et al., 2019b; Oroojlooy et al., 2020; Chenguang et al., 2021; Wu et al., 2021; Zhang et al., 2022b; Wei et al., 2019a; Wu et al., 2023). Transportation methods primarily focus on crafting efficient heuristic algorithms, dynamically adapting traffic signal configurations based on lane-level traffic conditions. However, these methods heavily depend on manual design and often exhibit limited effectiveness in complex scenarios. The emergence of deep neural networks (DNNs) (Mnih et al., 2015) led to the introduction of deep RL-based techniques to address this challenge (Abdulhai et al., 2003; El-Tantawy and Abdulhai, 2010). These approaches have exhibited remarkable performance across various traffic scenarios. Nevertheless, RL-based methods also present several drawbacks. Primarily, they may struggle with limited generalization ability, particularly when transferring to larger road networks or under highly uncommon scenarios (e.g., extreme high-traffic situations), as their training data only covers limited traffic situations. Additionally, RL-based methods lack interpretability due to the black-box nature of DNNs, which makes it hard to explain the rationale behind control actions under specific traffic conditions. This lack of transparency significantly hampers their feasibility for real-world deployment.
The recent emergence of Large Language Models (LLMs) has revolutionized several domains with their exceptional zero-shot learning and generalization capabilities, emulating human-like reasoning to solve complex tasks. For instance, AutoGPT (et al., 2023) proposes to break tasks into multiple sub-goals and iterate until the main task is completed. In the transportation domain, GPT-Driver (Mao et al., 2023) proposes to instruct GPT-3.5 to tackle motion planning tasks in autonomous driving. For traffic control, PromptGAT (Da et al., 2023) uses LLMs to generate human knowledge to help the DNN model understand long-tail cases (e.g., extreme weather) in TSC tasks, aiming to bridge the gap between real-world and simulations. Despite these advancements, existing research heavily relies on generalist LLMs (Wang et al., 2024) (e.g., ChatGPT) and primarily uses them as auxiliary tools (Da et al., 2023; Pang et al., 2024). However, generalist LLMs often lack specialized domain knowledge in TSC, potentially limiting their effectiveness in professional applications. The direct application of LLMs as TSC agents for human-like decision-making remains largely unexplored. Addressing these issues involves overcoming two major challenges.
First, LLMs are typically pre-trained on large-scale natural language corpora and rarely incorporate non-textual traffic data, such as sensor readings and GPS trajectories (Touvron et al., 2023; Chung et al., 2022). Despite their generalization capability across various tasks and domains, an inherent gap exists between real-time traffic data and linguistic understanding. The first challenge lies in enabling LLMs to comprehend real-time traffic dynamics and effectively interact with the traffic environment, which is critical for effective LLM-based TSC.
@@ 1. Introduction
To this end, we introduce LLMLight, a traffic signal control agent framework based on LLMs, as depicted in Figure 1. Specifically, we consider TSC as a partially observable Markov Game, where each agent, armed with an LLM, manages the traffic light at an intersection. At each signal-switching time step, the agent collects traffic conditions of the target intersection and verbalizes them into human-readable text as real-time observation. Additionally, we incorporate task descriptions enriched with commonsense knowledge about a control strategy to aid the LLM’s understanding of traffic management tasks. The combined real-time observation, task description, and control action space form a knowledgeable prompt that guides the LLM’s decision-making. Then, the agent leverages Chain-of-Thought (CoT) reasoning to determine the optimal traffic signal configuration for the subsequent time step.
@@
Furthermore, we construct a specialized LLM, LightGPT, to enhance the LLMLight framework. On the one hand, we propose imitation fine-tuning, enabling the specialized LLM to exploit high-quality control actions and underlying rationales derived from GPT-4. On the other hand, we introduce a policy refinement process that utilizes a well-trained critic model to evaluate and improve the agent’s actions. The optimized LightGPT produces more effective control policies and showcases remarkable generalization ability across diverse traffic scenarios in a more cost-effective manner than the leading LLMs, including GPT-4.
@@
We summarize the major contributions of this paper as follows: (1) We pioneer the application of advanced LLMs as control agents, enabling human-like decision-making for interpretable traffic signal control. To our knowledge, this is the first exploration of directly using LLMs as decision-making agents in TSC. (2) We introduce LLMLight, a tailored framework that engages various LLMs in effectively regulating traffic signals via knowledgeable instructions. (3) We construct LightGPT, a specialized LLM backbone optimized for TSC tasks. This significantly enhances the framework’s capability to derive effective control policies cost-effectively. (4) Extensive experiments on ten traffic datasets, along with evaluations by fifteen human experts, demonstrate superior effectiveness, generalization capabilities, and interpretability of LLMLight and LightGPT compared to nine baseline methods and ten advanced generalist LLMs. Additionally, these experiments reveal key insights into the potential of integrating LLMs into intelligent transportation systems.
@@ 2. Preliminary
Please refer to Figure 8 (See Appendix A.1) for more details of the most used setting of the intersection, lanes, and signal phases.
@@ 3.2.1. Observation Feature Construction
Subsequently, these features on every incoming lane (where the vehicle enters the intersection) are verbalized into human-readable text as the observation $o$ of the LLM.
@@ 3.2.2. Commonsense Augmented Prompt Construction
In addition to observation, we provide the LLM with comprehensive information essential in addressing traffic signal control, including traffic scenario description $d_{scene}$, task descriptions $d_{task}$, and control action space $A$. This enables the LLM to thoroughly understand the task, facilitating reasonable control actions. Furthermore, we integrate commonsense knowledge $d_{know}$ to mitigate generalist LLMs’ limitations in traffic management, a domain they are not explicitly trained for. Specifically, this knowledge dictates that the agent needs to prioritize lanes with extended queue lengths while paying less attention to vehicles distant from the intersection. Formally, we denote the composed prompt as $X=Prompt(o,d_{scene},d_{task},d_{know},A)$. A glimpse of the prompt template is shown in Figure 2(a), and a detailed example is provided in Table 11 (see Appendix A.12).
@@ 3.3. Backbone LLM Training
Fine-tuning has become a common routine to align LLM with diverse downstream applications (Chen et al., 2023a; Zeng et al., 2023) by augmenting the domain expertise in the specific realm. We propose a training procedure to optimize an LLM specifically tailored for traffic signal control, leading to the development of LightGPT. Figure 3 presents the training workflow. It primarily consists of three stages: (1) reasoning trajectory collection and filtering: it collects the CoT reasoning trajectory of GPT-4 for imitation fine-tuning, then we select the trajectories that most align with the long-term goal to ensure the data quality; (2) imitation fine-tuning: it utilizes the filtered reasoning trajectories and the corresponding actions taken by GPT-4 to train the backbone LLM; (3) critic-guided policy refinement: it further improves the decision-making process of the LLM by fine-tuning with the feedback of a well-trained critic model.
@@ 3.3.1. Trajectory Collection and Filtering
where $o_{t}$ and $a_{t}$ are the observation and action chosen at the signal-switching time step $t$, and $\gamma\in[0,1]$ is the discount factor. $R(o,a)$ is the reward function that provides feedback for executing action $a$ under observation $o$ (e.g., the negative of queue length). $Q(o,a)$ is the action-value function estimating the cumulative future reward earned after executing $a$. Actions estimated to yield higher future rewards suggest greater alignment with the long-term goal.
Subsequently, the trained action-value function is utilized as a critic to evaluate the decision-making of GPT-4. We selectively retain only reasoning trajectories where the selected actions are associated with the highest estimated future rewards, as:
@@ 3.3.2. Imitation Fine-tuning
where $P_{\pi_{\theta}}(y_{w}|X,Y_{<w})$ is the probability of being prompted with $X$ to generate the token $y_{w}$. Due to the enormous GPU memory usage of full model training, we adopt parameter-efficient fine-tuning (PEFT) techniques, specifically low-rank adaptation (LoRA) (Hu et al., 2022), in our training. Instead of adjusting all model weights, LoRA introduces low-rank matrices to model the weight updates, significantly reducing the number of parameters to be trained. This approach not only decreases computational and storage costs but also preserves the original model’s performance.
@@ 3.3.3. Critic-guided Policy Refinement
To further enhance the effectiveness of the LLM’s control policy, we propose a policy refinement procedure, which adjusts the LLM’s reasoning trajectory for directing better actions. Similarly, we use the pre-trained action-value function elaborated in Section 3.3.1 as the critic model to evaluate the actions derived by the backbone LLM. Higher evaluation scores correspond to greater estimated future rewards. Subsequently, we implement an alignment fine-tuning algorithm to adjust the reasoning trajectory, ultimately guiding the LLM towards actions that yield higher scores. Although previous studies (Ouyang et al., 2022; Carta et al., 2023) prove RL (e.g., PPO (Schulman et al., 2017)) is a potent tool for alignment, this method proved challenging to converge and yielded unsatisfactory results, as validated by (Wang et al., 2023a) and the experiment in Appendix A.9. Alternatively, we leverage a ranking loss (Wang et al., 2023a; Yuan et al., 2023) function to achieve this goal.
@@
where $p_{j*}=\min\limits_{q_{k}>q_{j}}p_{k}$ is the probability of the least favorable reasoning trajectory rated higher than $Y_{j}$, $\beta$ is a hyper-parameter, $e^{(p_{j}-p_{i})}$ is the alignment term to elevate the probability of generating trajectories with actions rated higher scores, and $e^{(2p_{j*}-2\beta-p_{i}-p_{j})}$ is the constraint term for prohibiting performance degradation.
It is crucial to be aware that while this refinement process is guided by a pre-trained critic model developed from RL, it does not imply that the backbone LLM learns exclusively from it. The alignment process is trained on reasoning trajectories sampled by the LLM’s policy, aiming to help the model identify trajectories with reasonable decisions that align closely with long-term goals. However, it is important to note that we do not provide any ground-truth labels in the training, thus the LLM is not bound to follow the critic’s predicted “best” action directly. Experiments in Subsection 4.7 demonstrate that RL models show performance degradation in varied traffic scenarios. In contrast, LLMLight equipped with LightGPT remains robust based on our proposed training procedure.
@@ 4.1. Experimental Settings
4.1.1. Datasets
Our experiments use seven real-world (Wei et al., 2019c) and three synthetic traffic flow datasets. Statistics for these datasets are detailed in Table 1. Visualizations of road networks are presented in Figure 18 (see Appendix A.11).
  • Jinan: This dataset is from Dongfeng sub-district, Jinan, China, featuring 12 intersections. It includes three traffic flow datasets recorded over three different periods. The area covered is 400 meters east-west and 800 meters north-south.
  • Hangzhou: Collected from Gudang sub-district, Hangzhou, China, this dataset features 16 intersections and includes two traffic flow datasets from different periods. The area spans 800 meters east-west and 600 meters north-south.
  • New York: Collected in the Upper East Side of Manhattan, New York, USA, using taxi trip data, this dataset covers 196 intersections with lanes spanning 300 meters. It includes two large-scale traffic flow datasets from different periods to evaluate the scalability of different methods.
  • Extreme High-traffic: Synthesized based on real-world traffic in Jinan and Hangzhou, these datasets feature approximately four times more vehicles per five minutes than the original datasets to simulate high-traffic scenarios.
  • 24-hour Cycle: Modeled on Jinan’s actual traffic flows, this dataset replicates realistic daily traffic fluctuations over a 24-hour period, including rush hours (Adeyemi et al., 2021) from 6:00 am to 10:00 am and 3:00 pm to 8:00 pm.
@@ 4.1.1. Datasets
4.1.2. Data Anonymization
The datasets we used do not include any identifiable driver information, such as names or contact details. These datasets are open-sourced and publicly available at (Zhang et al., 2023a).
@@ 4.1.3. Environment Settings
We conduct experiments on CityFlow, an open-source simulator widely used for benchmarking traffic signal control policies. CityFlow supports large-scale, city-wide traffic simulations and features an interactive render for monitoring. It provides extensive APIs to access traffic state features and execute control actions. In our experiments, each intersection connects four lane sections: East, West, North, and South. We utilize four signal phases as corresponding control actions: ETWT (go-through from east and west), ELWL (left-turn from east and west), NTST (go-through from north and south), and NLSL (left-turn from north and south). Right-turn movements are always allowed in the simulation. The green signal phase duration is set to thirty seconds. Following existing studies (Zhang et al., 2022b; Wei et al., 2019b), a three-second yellow signal and a two-second all-red time follow each green signal to prepare the transition. The simulation spans twenty-four hours for the 24-hour cycle dataset, while others span one hour.